OpenStudy (anonymous):

If you have a group of 6 men and 4 women, how many groups of 5 can you make that have at least one woman?

OpenStudy (shubhamsrg):

there will be 2 groups with 5 members each no. of woman in each group can be 1 and 3 2 and 2 3 and 1 So possible combinations are : C(4,1)*C(6,4) + C(4,2)*C(6,3) + C(4,3)*C(6,2)

OpenStudy (anonymous):

Why are you adding them up to 4? It's a group of 5 members

OpenStudy (shubhamsrg):

I did not explain well. Leme try this way. in 1 group,either there will be 1 woman and 4 men, or 2 women, 3 men or 3 women , 2 men there can not be 4 women in a group as that will leave other group as 0 women. Hence possibilities are C(4,1) * C(6,4) or C(4,2) * C(6,3) or C(4,3) * C(6,2)

OpenStudy (shubhamsrg):

We shall add all them for the final result.

OpenStudy (shubhamsrg):

Oh wait, maybe I misunderstood the ques.

OpenStudy (shubhamsrg):

There can be 4 women in a group since we are forming just 1 group.hmm So we will add C(4,4)*C(6,1) as well.

OpenStudy (anonymous):

I'll type it as it shows up on the homework I just changed up the values a bit A committee of 5 people must be chosen from a group of 6 men and 4 women. If the committee is required to have at least 1 woman, how many different committee choices are possible?

OpenStudy (anonymous):

one way to go is to compute the chance (x) to have a group without women, and then do 1-x.

OpenStudy (anonymous):

But if I'm calculating the number of possible combinations, how would I use chance?

OpenStudy (anonymous):

correct, I read the question too fast :-(

OpenStudy (anonymous):

Lol it's ok

OpenStudy (agent0smith):

I think @ataly might have been on the right track... find the total number of possible combinations, and then subtract #groups without women.

OpenStudy (agent0smith):

Total possible combinations = 10C5 # without women = 6C5 10C5 - 6C5 = 252 - 6 = 246 (not 100% on this, haven't done combinations in a while)

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (agent0smith):

No prob. Hopefully @shubhamsrg or someone will check that. I think it's right though... total groups is: from 10 people choose 5 people, groups without women are: from 6 men choose 5 men.

OpenStudy (agent0smith):

The long way would be like so: 6C4*4C1 + 6C3*4C2 + 6C2*4C3 + 6C1*4C4 = 246 So both ways get the same answer... which is usually a good sign!